# Align Your Rhythm: Generating Highly Aligned Dance Poses with Gating-Enhanced Rhythm-Aware Feature Representation

### v1

@@ Abstract
Automatically generating natural, diverse and rhythmic human dance movements driven by music is vital for virtual reality and film industries. However, generating dance that naturally follows music remains a challenge, as existing methods lack proper beat alignment and exhibit unnatural motion dynamics. In this paper, we propose Danceba, a novel framework that leverages gating mechanism to enhance rhythm-aware feature representation for music-driven dance generation, which achieves highly aligned dance poses with enhanced rhythmic sensitivity. Specifically, we introduce Phase-Based Rhythm Extraction (PRE) to precisely extract rhythmic information from musical phase data, capitalizing on the intrinsic periodicity and temporal structures of music. Additionally, we propose Temporal-Gated Causal Attention (TGCA) to focus on global rhythmic features, ensuring that dance movements closely follow the musical rhythm. We also introduce Parallel Mamba Motion Modeling (PMMM) architecture to separately model upper and lower body motions along with musical features, thereby improving the naturalness and diversity of generated dance movements. Extensive experiments confirm that Danceba outperforms state-of-the-art methods, achieving significantly better rhythmic alignment and motion diversity. Project page: https://danceba.github.io/.
@@ 1 Introduction
Music-driven dance generation aims to automatically synthesize realistic and expressive 3D dance movements synchronized precisely with a given musical sequence, enabling natural animation of 3D character models. Due to its broad applications in gaming, virtual reality, film production, and interactive entertainment [11, 33, 29], this task has attracted considerable attention. However, generating dances that accurately reflect musical rhythm while preserving natural and diverse movement dynamics remains a significant challenge, as it requires effectively capturing and utilizing complex temporal relationships between music and motion.
@@
To further tackle the rhythm alignment issue, Beat-It [37] explicitly introduces beat synchronization via a beat distance predictor. However, such explicit synchronization approach impose rigid control constraints and strong regularization during generation, significantly limiting the diversity and expressiveness of generated dances [23]. Thus, a more balanced approach capable of achieving precise rhythmic alignment without sacrificing motion naturalness and diversity is highly desirable.
To this end, we propose a novel framework equipped with a gating-enhanced rhythm-aware feature representation for music-driven dance generation, i.e., Danceba, which achieves highly aligned dance poses with enhanced rhythmic sensitivity. Specifically, Danceba introduces Phase-Based Rhythm Extraction (PRE), a rhythm-aware feature extraction module that explicitly captures rhythmic characteristics from musical phase information. Leveraging the intrinsic temporal phase information intrinsic to rhythm perception, PRE effectively separates rhythmic from semantic music features, significantly improving the rhythmic alignment between generated dance sequences and musical inputs.
To further enhance beat-dance alignment [^1] In this paper, the terms “beat” and “rhythm” are used interchangeably in music processing following Diffdance [23] and Bailando++ [27]. using the rhythm-aware features extracted by PRE, it is essential to establish a stronger contextual association between musical rhythm and dance movements. To achieve this, we present a Temporal-Gated Causal Attention (TGCA) mechanism, enhanced by a gating strategy to explicitly prioritize and leverage global rhythmic structures within cross-conditional causal attention. By integrating gating mechanism, TGCA effectively strengthens the contextual connection between music rhythm and generated movements, ensuring that generated dance sequences accurately reflect the underlying musical rhythm and exhibit greater coherence and expressive diversity.
@@ 2.1 Music-Driven Dance Generation
Music-driven dance generation focuses on automatically synthesizing 3D dance movements from musical input, creating temporally coherent pose sequences for animating 3D character models. Early approaches, including [13, 39, 10, 14], investigated diverse methodologies for this task. Li et al. [13] utilized a standard Transformer [30] with a late-fusion module to predict subsequent dance steps. The FACT framework [14] introduced the AIST++ dataset and employed a cross-modal transformer with full attention mechanisms for music-to-dance generation. However, these initial methods neglect the rhythm features of music and struggled to effectively integrate musical and motion features, leading to suboptimal generation quality.
@@ 2.2 Mamba in Human Motion Generation
Figure 1: The overall framework of Danceba consists of three core modules: Phase-based Rhythm Extraction (PRE), Temporal-Gated Causal Attention (TGCA), and Parallel Mamba Motion Modeling (PMMM). PRE precisely extracts rhythm-aware features from musical phase information, providing accurate rhythmic signals. These rhythm-enhanced music features are fused with pose embeddings (upper-body and lower-body poses) and processed through the TGCA module, which utilizes gating mechanism to reinforce rhythmic sensitivity and align dance movements accurately with music beats. Parallel Mamba Motion Modeling separately models upper and lower body motion sequences, effectively capturing distinct dance dynamics aligned with rhythm-aware music feature, thus significantly enhancing the naturalness, diversity, and temporal coherence of the generated dance movements.
@@ 3.1 Preliminary
At each time step $t$, MotionGPT [26] estimates the probabilities of pose codes $\mathrm{z}_{i}\in\mathbb{Z}$ and select the one with the most probable pose codes as the predicted upper pose $\hat{p}_{t}^{u}$ and predicted lower pose $\hat{p}_{t}^{l}$.
@@ 3.2 Phase-Based Rhythm Feature Extraction
$\boldsymbol{\varphi}^{\prime}$ $=\operatorname{Angle}\left(\mathbf{S}_{m}\right),$
@@
where $\boldsymbol{\varphi}\in\mathbb{C}^{T^{\prime}\times D_{\varphi}}$ has the same temporal length $T^{\prime}$, addressing the potential temporal missalignments and preserving critical rhythmic cues.
@@ 3.3 Global Beat Attention via Temporal Gating
Leveraging these rhythm-rich features, we next investigate how to integrate them into the alignment process between music and dance. Howover, prior methods [26, 38, 27, 9, 31] relying on cross-conditional causal attention mechanism (denoted as $\operatorname{C^{3}Attention}$) exhibit a critical misalignment issue. As shown in Figure 3 (a), we visualize the cross-conditional causal attention mechanism and observe that the next predicted token lacks direct control signals from historical tokens, failing to establish a stable global beat attention. This misalignment leads to error accumulation, which we identify as a key factor behind the unnatural dance generation in prior work.
Figure 3: Heat map of attention. The two displayed images show the results learned by Cross-Condional Causal Attention (a) and Temporal-Gated Causal Attention (b) under the same piece of music. The brighter the color, the higher the attention; the darker the color, the lower the attention. TGCA the next token has a clear control signal, preventing overly random dance movements and achieving enhanced rhythm-aware feature representation.
@@ 3.4 Mamba-Based Parallel Motion Modeling
After TGCA enhances dance movements in $\mathbf{X}_{attn}$ with global rhythmic perception, we looking into dance modeling for generation. Dance generation faces two key challenges: (1) Dance is a complex spatio-temporal sequence encoding trajectory, posture, and velocity over time, requiring a model capable of capturing both local transitions and global rhythmic variations. (2) The upper and lower body exhibit distinct movement patterns and speeds, necessitating separate processing to maintain coherence and diversity. To address these issues, we propose a parallel Mamba motion modeling architecture that separately models music, upper-body, and lower-body movements while leveraging state space modeling to ensure fluid and diverse motion generation.
@@
The gating mechanism dynamically regulates the frequency, speed, and amplitude of movements in response to musical beats, ensuring that dance sequences remain synchronized with the music. Finally, three parallel Mamba architectures model upper- and lower-body dance sequences separately, therefore ensures the generated dance movements are expressive, diverse, and temporally coherent.
@@ 4.1 Experimental Setup
Dataset: Following existing methods [26, 23, 31], we conduct all experiments on AIST++ dataset [14], which is the most widely used dataset for music-driven dance generation. AIST++ consists of 991 high-quality 3D pose sequences recorded at 60 FPS in skinned multi-person linear (SMPL) format [18], with 951 sequences designated for training and 40 reserved for evaluation. Following [26, 23, 31], we generate 40 pieces of dance sequences on the test set of AIST++ dataset, and sample the generated dance sequence with length of 20 seconds for performance evaluation.
@@ 4.2 Comparison with State-of-the-Art Methods
To validate the superiority of our method, we conduct comparative experiments with the state-of-the-art methods [13, 10, 39, 14, 26, 28, 23, 27, 31, 16], as well as the ground truth. Following [26, 23, 31], we generate 40 pieces of dance sequences on the AIST++ test set, and sample the generated dance sequence with the length of 20 seconds to compute the metrics mentioned above. Among the compared methods, “Li et al.” [13], DanceNet [39] and FACT [14] use the results from the AIST++ benchmark [14], while the results of DanceRevolution [10] are derived from its reproduction as presented in Bailando [26]. EDGE [28] utilizes the reproduced results from the Lodge [16]. The Bailando, DiffDance [23] and Lodge methods use the experimental results from their respective publications.
@@
Motion Quality Metrics: As illustrated in Table 1, our method outperforms existing methods in overall quality performance. To be specific, our method improves on $\text{FID}_{k}$ by 48.68% (11.67 vs. 22.74), while exhibits competitive edge of $\text{FID}_{g}$. Further analysis of the employed quality metrics reveals distinct roles for the two metrics types. The metric $\text{FID}_{k}$, defined based on velocity and energy, primarily captures the kinetic characteristics of the generated dance, while $\text{FID}_{g}$, derived from multiple manually designed motion templates, emphasizes the geometric structural properties of the dance. Thus, our method demonstrates superior performance in modeling the kinetic aspects of sequential motions, as evidenced by its exceptional $\text{FID}_{k}$ scores. Although it exhibits a modest gap in $\text{FID}_{g}$ compared to the state-of-the-art Bailando++ [27], this difference is acceptable given its focus on kinetic fidelity. Despite this minor trade-off, our method outperforms existing approaches overall, achieving a robust balance of efficiency and quality across a broader range of metrics.
Motion Diversity Metrics: Regarding diversity evaluation, Danceba achieves the best diversity performance, which improves $\text{Div}_{k}$ by 7.0% (8.52 vs. 7.96) and even slightly outperforms the Ground Truth (8.52 vs. 8.19). Meanwhile, our method improves $\text{Div}_{g}$ by 16.3% (7.55 vs. 6.49) and also slightly surpasses the Ground Truth (7.55 vs. 7.45), as illustrate in Table 1. For a visual comparison, we also present a comparison with the state-of-the-art method Bailando, as illustrate in Figure 5. It can be observed that Bailando exhibits relatively small movement amplitudes, with some dance movements missing and the actions appearing stiff. In contrast, our method shows larger movement amplitudes and movement diversity.
@@
Figure 6: Comparison with Bailando++ [27]. Danceba achieves better rhythmic alignment, enhancing synchronization between dance movements and musical rhythm.
@@ 4.3 Ablation Studies
We conduct ablation studies to evaluate the contributions of the key modules (i.e., PRE, TGCA, and PMMM) in Danceba, where “w/o” denotes the removal of a specific module. The results are presented in Table 2. Additionally, we assess our method using a single Mamba structure instead of the parallel Mamba structure for motion modeling, referred to as Danceba-Single, with results shown in Table 3.
@@
Table 2: Ablation study of three key modules (i.e., PRE, TGCA, and PMMM), and “w/o” indicates Danceba without specified module. Visual comparisons with Bailando can be found in the supplementary video.
@@
Phase-Based Rhythm Extraction: To validate the effectiveness of PRE, we replace the rhythm feature extraction module with a learnable linear layer as used in previous works [26, 38, 27, 9, 31]. Consequently, the motion quality (i.e., FID), motion diversity (i.e., Div), and beat alignment score (i.e., BAS) decrease by 6.09 (41.91%), 1.14 (14.86%), and 0.0305 (10.96%), respectively, as illustrated in Table 2. Specifically, this performance drop highlights the contribution of RPE, which effectively captures the critical rhythm features of music, a crucial aspect that previous methods have largely overlooked.
@@ 5 Limitations
Although Danceba achieves superior dance generation quality compared to state-of-the-art methods, it has two key limitations: (1) Music feature encoding: Danceba uses a simple linear layer instead of a pre-trained audio model (e.g., Jukebox [2], MERT [17], or CLAP [4]). While computationally efficient, this design may fail to capture nuanced musical structures, affecting rhythm-motion synchronization. Future work could explore pre-trained music encoders to enhance feature extraction. (2) 3D pose quantization: Danceba adopts the Pose VQ-VAE framework from Bailando [26], which has a limited encoding space, restricting motion diversity. This trade-off between quantization efficiency and expressiveness may hinder fine-grained motion details. Exploring hierarchical quantization or adaptive codebook learning could further improve motion quality.

### v2

@@ Abstract
Automatically generating natural, diverse, and rhythmic human dance movements driven by music is vital for virtual reality and film industries. However, generating dance that naturally follows music remains a challenge, as existing methods lack proper beat alignment and exhibit unnatural motion dynamics. In this paper, we propose Danceba, a novel framework that leverages gating mechanism to enhance rhythm-aware feature representation for music-driven dance generation, which achieves highly aligned dance poses with enhanced rhythmic sensitivity. Specifically, we introduce Phase-Based Rhythm Extraction (PRE) to precisely extract rhythmic information from musical phase data, capitalizing on the intrinsic periodicity and temporal structures of music. Additionally, we propose Temporal-Gated Causal Attention (TGCA) to focus on global rhythmic features, ensuring that dance movements closely follow the musical rhythm. We also introduce Parallel Mamba Motion Modeling (PMMM) architecture to separately model upper and lower body motions along with musical features, thereby improving the naturalness and diversity of generated dance movements. Extensive experiments confirm that Danceba outperforms state-of-the-art methods, achieving improved rhythmic alignment and motion diversity. Project page: https://danceba.github.io/.
@@ 1 Introduction
Music-driven dance generation aims to automatically synthesize realistic and expressive 3D dance movements synchronized precisely with a given musical sequence, enabling natural animation of 3D character models. Due to its broad applications in gaming, virtual reality, film production, and interactive entertainment [43, 14, 38, 34], this task has attracted considerable attention. However, generating dances that accurately reflect musical rhythm while preserving natural and diverse movement dynamics remains a significant challenge, as it requires effectively capturing and utilizing complex temporal relationships between music and motion.
@@
To further tackle the rhythm alignment issue, Beat-It [42] explicitly introduces beat synchronization via a beat distance predictor. However, such an explicit synchronization approach imposes rigid control constraints and strong regularization during generation, significantly limiting the diversity and expressiveness of generated dances [27]. Thus, a more balanced approach capable of achieving precise rhythmic alignment without sacrificing motion naturalness and diversity is highly desirable.
To this end, we propose a novel framework equipped with a gating-enhanced rhythm-aware feature representation for music-driven dance generation, i.e., Danceba, which achieves highly aligned dance poses with enhanced rhythmic sensitivity. Specifically, Danceba introduces Phase-Based Rhythm Extraction (PRE), a rhythm-aware feature extraction module that explicitly captures rhythmic characteristics from musical phase information. Leveraging the temporal phase information intrinsic to rhythm perception, PRE effectively separates rhythmic from semantic music features, significantly improving the rhythmic alignment between generated dance sequences and musical inputs.
To further enhance beat-dance alignment [^1] Note that, the terms “beat” and “rhythm” are used interchangeably in music processing following DiffDance [27] and Bailando++ [31]. using the rhythm-aware features extracted by PRE, it is essential to establish a stronger contextual association between musical rhythm and dance movements. To achieve this, we present a Temporal-Gated Causal Attention (TGCA) mechanism, enhanced by a gating strategy to explicitly prioritize and leverage global rhythmic structures within cross-conditional causal attention. By integrating gating mechanism, TGCA effectively strengthens the contextual connection between music rhythm and generated movements, ensuring that generated dance sequences accurately reflect the underlying musical rhythm and exhibit greater coherence and expressive diversity.
@@ 2.1 Music-Driven Dance Generation
Music-driven dance generation focuses on automatically synthesizing 3D dance movements from musical input, creating temporally coherent pose sequences for animating 3D character models. It has attracted increasing attention in both the computer vision and virtual avatar communities [13, 1] due to its applications in entertainment, gaming, and virtual reality. Early approaches, including [10, 15, 32, 17, 45, 12, 18], investigated diverse methodologies for this task. The study in [17] utilized a standard Transformer [35] with a late-fusion module to predict subsequent dance steps. The FACT framework [18] introduced the AIST++ dataset and employed a cross-modal Transformer with full attention mechanisms for music-to-dance generation. However, these initial methods neglect the rhythm features of music and struggle to effectively integrate musical and motion features, leading to suboptimal generation quality.
@@
Figure 1: The overall framework of Danceba consists of three core modules: Phase-based Rhythm Extraction (PRE), Temporal-Gated Causal Attention (TGCA), and Parallel Mamba Motion Modeling (PMMM). PRE precisely extracts rhythm-aware features from musical phase information, providing accurate rhythmic signals. These rhythm-enhanced music features are fused with pose embeddings (upper-body and lower-body poses) and processed through the TGCA module, which utilizes gating mechanism to reinforce rhythmic sensitivity and align dance movements accurately with music beats. Parallel Mamba Motion Modeling separately models upper and lower body motion sequences, effectively capturing distinct dance dynamics aligned with rhythm-aware music features, thus significantly enhancing the naturalness, diversity, and temporal coherence of the generated dance movements.
@@ 3.1 Preliminary
At each time step $t$, MotionGPT [30] estimates the probabilities of pose codes $\mathrm{z}_{i}\in\mathbb{Z}$ and selects the one with the most probable pose codes as the predicted upper pose $\hat{p}_{t}^{u}$ and predicted lower pose $\hat{p}_{t}^{l}$.
@@ 3.2 Phase-Based Rhythm Feature Extraction
$\boldsymbol{\varphi}$ $=\operatorname{Angle}\left(\mathbf{S}_{m}\right),$
@@
where $\boldsymbol{\varphi}\in\mathbb{C}^{T^{\prime}\times D_{\varphi}}$ has the same temporal length $T^{\prime}$, addressing the potential temporal misalignment and preserving critical rhythmic cues.
@@ 3.3 Global Beat Attention via Temporal Gating
Leveraging these rhythm-rich features, we next investigate how to integrate them into the alignment process between music and dance. However, prior methods [30, 44, 31, 11, 36] relying on cross-conditional causal attention mechanism (denoted as $\operatorname{C^{3}Attention}$) exhibit a critical misalignment issue. As shown in Figure 3 (a), we visualize the cross-conditional causal attention mechanism and observe that the next predicted token lacks direct control signals from historical tokens, failing to establish a stable global beat attention. This misalignment leads to error accumulation, which we identify as a key factor behind the unnatural dance generation in prior work.
Figure 3: Heat map of attention. The two displayed images show the results learned by Cross-Conditional Causal Attention (a) and Temporal-Gated Causal Attention (b) under the same piece of music. The brighter the color, the higher the attention; the darker the color, the lower the attention. TGCA the next token has a clear control signal, preventing overly random dance movements and achieving enhanced rhythm-aware feature representation.
@@ 3.4 Mamba-Based Parallel Motion Modeling
After TGCA enhances dance movements in $\mathbf{X}_{attn}$ with global rhythmic perception, we looking into dance modeling for generation. Dance generation faces two key challenges: (1) Dance is a complex spatial-temporal sequence encoding trajectory, posture, and velocity over time, requiring a model capable of capturing both local transitions and global rhythmic variations. (2) The upper and lower body exhibit distinct movement patterns and speeds, necessitating separate processing to maintain coherence and diversity. To address these issues, we propose a parallel Mamba motion modeling architecture that separately models music, upper-body, and lower-body movements while leveraging state space modeling to ensure fluid and diverse motion generation.
@@
The gating mechanism dynamically regulates the frequency, speed, and amplitude of movements in response to musical beats, ensuring that dance sequences remain synchronized with the music. Finally, three parallel Mamba architectures model upper- and lower-body dance sequences separately, therefore ensure the generated dance movements are expressive, diverse, and temporally coherent.
@@ 4.1 Experimental Setup
Dataset: We follow prior works [30, 27, 36] and perform all experiments on the AIST++ dataset [18], the most widely used benchmark for music-driven dance generation that avoids observable biases (e.g., offensive postures). AIST++ consists of 991 high-quality 3D pose sequences recorded at 60 FPS in skinned multi-person linear (SMPL) format [22], with 951 sequences designated for training and 40 reserved for evaluation. Following [30, 27, 36], we generate 40 pieces of dance sequences on the test set of AIST++ dataset, and sample the generated dance sequence with length of 20 seconds for performance evaluation.
@@ 4.2 Comparison with State-of-the-Art Methods
To validate the superiority of our method, we conduct comparative experiments with the state-of-the-art methods [17, 12, 45, 18, 30, 33, 27, 31, 36, 20], as well as the ground truth. Following [30, 27, 36], we generate 40 pieces of dance sequences on the AIST++ test set, and sample the generated dance sequence with the length of 20 seconds to compute the metrics mentioned above. Among the compared methods, “Li et al.[17], DanceNet [45] and FACT [18] use the results from the AIST++ benchmark [18], while the results of DanceRevolution [12] are derived from its reproduction as presented in Bailando [30]. EDGE [33] utilizes the reproduced results from the Lodge [20]. The Bailando, DiffDance [27] and Lodge methods use the experimental results from their respective publications.
@@
Motion Quality Metrics: As illustrated in Table 1, our method outperforms existing methods in overall quality performance. To be specific, our method improves on $\text{FID}_{k}$ by 48.68% (11.67 vs. 22.74), while exhibits competitive edge of $\text{FID}_{g}$. Further analysis of the employed quality metrics reveals distinct roles for the two metrics types. The metric $\text{FID}_{k}$, defined based on velocity and energy, primarily captures the kinetic characteristics of the generated dance, while $\text{FID}_{g}$, derived from multiple manually designed motion templates, emphasizes the geometric structural properties of the dance. Thus, our method demonstrates superior performance in modeling the kinetic aspects of sequential motions, as evidenced by its exceptional $\text{FID}_{k}$ scores. Although it exhibits a modest gap in $\text{FID}_{g}$ (11.90 vs. 11.58) compared to the state-of-the-art Bailando++ [31], it achieves a 19.0% higher $\text{Div}_{g}$ (7.55 vs. 6.34), indicating improved motion diversity with comparable generative quality. Despite this minor trade-off, our method outperforms existing approaches overall, achieving a robust balance of efficiency and quality across a broader range of metrics.
Motion Diversity Metrics: Regarding diversity evaluation, Danceba achieves the best diversity performance, which improves $\text{Div}_{k}$ by 7.0% (8.52 vs. 7.96) and even slightly outperforms the Ground Truth (8.52 vs. 8.19). Meanwhile, our method improves $\text{Div}_{g}$ by 16.3% (7.55 vs. 6.49) and also slightly surpasses the Ground Truth (7.55 vs. 7.45), as illustrate in Table 1. For a visual comparison, we also present a comparison with the state-of-the-art method Bailando, as illustrated in Figure 5. It can be observed that Bailando exhibits relatively small movement amplitudes, with some dance movements missing and the actions appearing stiff. In contrast, our method shows larger movement amplitudes and movement diversity.
@@
Figure 6: Beats alignment visualization, where the horizontal axis shows frame indices of beat events. Comparing Danceba with Bailando++ [31], we can find that the distance between motion beats and music beats generated by our method is smaller. This indicating that Danceba performs better in terms of rhythmic alignment.
@@ 4.3 Ablation Studies
We conduct ablation studies to evaluate the contributions of the key modules (i.e., PRE, TGCA, and PMMM) in Danceba, where “w/o" denotes the removal of a specific module. “w/o PMMM” refers to the setting where PMMM is replaced with TGCA layers to ensure fairness. The results are presented in Table 2. Additionally, we assess our method using a single Mamba structure instead of the parallel Mamba structure for motion modeling, referred to as Danceba-Single, with results shown in Table 3.
@@
Table 2: Ablation study of three key modules (i.e., PRE, TGCA, and PMMM), and “w/o" indicates Danceba without specified module. Visual comparisons with Bailando can be found in the supplementary video.
@@
Phase-Based Rhythm Extraction: To validate the effectiveness of PRE, we replace the rhythm feature extraction module with a learnable linear layer as used in previous works [30, 44, 31, 11, 36]. Consequently, the motion quality (i.e., FID), motion diversity (i.e., Div), and beat alignment score (i.e., BAS) decrease by 6.09 (41.91%), 1.14 (14.86%), and 0.0305 (10.96%), respectively, as illustrated in Table 2. Specifically, this performance drop highlights the contribution of PPE, which effectively captures the critical rhythm features of music, a crucial aspect that previous methods have largely overlooked.
@@ 5 Limitations
Although Danceba achieves superior dance generation quality compared to state-of-the-art methods, it has two key limitations: (1) Music feature encoding: Danceba uses a simple linear layer instead of a pre-trained audio model (e.g., Jukebox [3], MERT [21], or CLAP [5]). While computationally efficient, this design may fail to capture nuanced musical structures, affecting rhythm-motion synchronization. Future work could explore pre-trained music encoders to enhance feature extraction. (2) 3D pose quantization: Danceba adopts the Pose VQ-VAE framework from Bailando [30] for fair comparison with prior SOTA methods [30, 31, 36]. However, it has a limited encoding space, which may limit motion diversity. Such a trade-off between quantization efficiency and expressiveness could hinder the preservation of fine-grained motion details. Exploring hierarchical quantization or adaptive codebook learning could further improve motion quality.